3.9GHz status
Monday, 4 August 2008 -
09:00
Monday, 4 August 2008
09:00
Cavity No. 6 vertical test results
Cavity No. 6 vertical test results
09:00 - 09:15
Room: The Hermitage
09:15
A0 update -- high pressure rinse system, North cave vacuum system, etc.
A0 update -- high pressure rinse system, North cave vacuum system, etc.
09:15 - 09:35
Room: The Hermitage
09:35
3rd harmonic helium supply weldments
3rd harmonic helium supply weldments
09:35 - 09:50
Room: The Hermitage
09:50
Repair of cavity No. 2
Repair of cavity No. 2
09:50 - 10:05
Room: The Hermitage
10:05
Other business
Other business
10:05 - 10:25
Room: The Hermitage